Solve the system using substitution.
y = 2x + 10
x^2 + y = 13

Answer by mananth(16949)

y = 2x + 10
x^2 + y = 13
x^2+(2x+10)=13
x^2+2x+10-13=0
x^2+2x-3=0
x^2-3x+x-3=0
x(x-3)+1(x-3)=0
(x-3)(x+1)=0
x=3,-1
y=16,8
(3,16) (-1,8)